# DynamicID: Zero-Shot Multi-ID Image Personalization with Flexible Facial Editability

### v2

@@ Abstract
Recent advances in text-to-image generation have driven interest in generating personalized human images that depict specific identities from reference images. Although existing methods achieve high-fidelity identity preservation, they are generally limited to single-ID scenarios and offer insufficient facial editability. We present DynamicID, a tuning-free framework that inherently facilitates both single-ID and multi-ID personalized generation with high fidelity and flexible facial editability. Our key innovations include: 1) Semantic-Activated Attention (SAA), which employs query-level activation gating to minimize disruption to the base model when injecting ID features and achieve multi-ID personalization without requiring multi-ID samples during training. 2) Identity-Motion Reconfigurator (IMR), which applies feature-space manipulation to effectively disentangle and reconfigure facial motion and identity features, supporting flexible facial editing. 3) a task-decoupled training paradigm that reduces data dependency, together with VariFace-10k, a curated dataset of 10,000 unique individuals, each represented by 35 distinct facial images. Experimental results demonstrate that DynamicID outperforms state-of-the-art methods in identity fidelity, facial editability, and multi-ID personalization capability. Our code is available at https://github.com/ByteCat-bot/DynamicID.
@@ 1 Introduction
In recent years, the field of text-to-image synthesis has seen remarkable progress, largely due to advancements in diffusion-based models [48, 59, 55, 54, 36]. This progress has spurred interest in personalized human image generation, a domain with extensive applications such as AI-generated portraits, image animation, and e-commerce advertising. The primary goal is to generate images that maintain consistent identities from reference images while seamlessly incorporating additional prompts.
Previous research in personalized human image generation has achieved substantial progress through various approaches [1, 52, 8, 4, 6, 14, 38, 74]. Early methodologies [57, 16, 58, 35] demonstrated exceptional high-fidelity outcomes but required subject-specific retraining, imposing significant computational burdens and temporal constraints. To address these limitations, subsequent tuning-free paradigms [42, 71, 82, 73, 28, 79] employed pre-trained face encoders coupled with feature injection modules. These frameworks extract facial features through the encoder and subsequently integrate them into the generation process via the injection module. While these parameter-efficient methods eliminate retraining requirements and improve practical deployment capabilities, two critical limitations persist: Constraints in Multi-ID Generation: In scenarios involving multiple identities, numerous approaches struggle with the challenge of identity blending [78]. While various mitigation strategies have emerged, these methods are built upon frameworks tailored for the single-ID contexts, which inherently compromises model performance. Insufficient Facial Attribute Editability: Current methodologies lack explicit training models to disentangle identity features (e.g., facial structure and skin texture) from motion features (e.g., expression and orientation). This intrinsic entanglement within facial latent representations fundamentally restricts the flexible manipulation of facial attributes while preserving identity fidelity.
@@ 2.1 Text-to-Image Models
The field of text-to-image generation has experienced significant advancements, driven by innovative techniques such as Generative Adversarial Networks (GANs) [30, 77], auto-regressive models [54, 66], and diffusion models [25, 55]. Among these techniques, diffusion models have emerged as the predominant approach, leveraging diffusion processes to iteratively transform noise into coherent images that align with textual descriptions [51, 48, 49, 65, 15]. Our methodology is built upon the stable diffusion model, which is in line with current trends.
@@ 2.2 Personalized Human Image Generation
Tuning-free single-ID personalization: A vast amount of tuning-free studies have concentrated on the single-ID contexts [5, 10, 9, 29, 44, 68], which typically utilize a face encoder to extract facial features from reference images and an injection module to integrate these features into the generative process. Current approaches can be systematically categorized into three distinct types based on the injection mechanism: text-based methods [67, 63, 7, 72], network-based methods [21, 82], and attention-based methods [43, 76, 17, 20, 75]. To illustrate, PhotoMaker [41] extracted facial features and merged them with the corresponding text token within the text embedding space, achieving semantic consistency though compromising identity fidelity. Conversely, InstantID [71] adapts ControlNet [81] architecture for high-fidelity generation but lacks facial editing granularity. Improving upon this, CapHuman [42] enhances facial editability via 3D Morphable Face Model [39] integration at the cost of increased computational complexity. As the pioneering attention-based method, IP-Adapter [80] introduces a novel decoupled cross-attention mechanism for facial identity integration, achieving superior fidelity while sharing InstantID’s limitation in granular facial control. Inspired by this, the $W_{+}$ adapter [40] leveraged the StyleGAN [32] encoder to extract features for expression control. However, using this encoder reduced fidelity and limited control over facial orientation. Moreover, these attention-based methods disrupted the model’s original behavior during the generative process due to the coarse-grained nature of cross-attention integration.
@@ 4.3 Ablation study
5 Limitations
Although our plug-and-play framework fundamentally supports all text-to-image models, this work primarily uses Stable Diffusion v1.5 [55], which limits performance when generating faces near image edges [69] or handling complex prompts. Future work will explore scaling via larger datasets, broader model adaptation, and version releases.

### v3

@@ Abstract
Recent advances in text-to-image generation have driven interest in generating personalized human images that depict specific identities from reference images. Although existing methods achieve high-fidelity identity preservation, they are generally limited to single-ID scenarios and offer insufficient facial editability. We present DynamicID, a tuning-free framework that inherently facilitates both single-ID and multi-ID personalized generation with high fidelity and flexible facial editability. Our key innovations include: 1) Semantic-Activated Attention (SAA), which employs query-level activation gating to minimize disruption to the base model when injecting ID features and achieve multi-ID personalization without requiring multi-ID samples during training. 2) Identity-Motion Reconfigurator (IMR), which applies feature-space manipulation to effectively disentangle and reconfigure facial motion and identity features, supporting flexible facial editing. 3) a task-decoupled training paradigm that reduces data dependency, together with VariFace-10k, a curated dataset of 10k unique individuals, each represented by 35 distinct facial images. Experimental results demonstrate that DynamicID outperforms state-of-the-art methods in identity fidelity, facial editability, and multi-ID personalization capability.
@@ 1 Introduction
In recent years, the field of text-to-image synthesis has seen remarkable progress, largely due to advancements in diffusion-based models [57, 73, 69, 68, 38, 45, 65, 106, 105]. This progress has spurred interest in personalized human image generation, a domain with extensive applications such as AI-generated portraits, image animation, and e-commerce advertising. The primary goal is to generate images that maintain consistent identities from reference images while seamlessly incorporating additional prompts.
Previous research in personalized human image generation has achieved substantial progress through various approaches [1, 66, 9, 4, 6, 16, 40, 92, 41, 52, 84]. Early methodologies [71, 18, 72, 37] demonstrated exceptional high-fidelity outcomes but required subject-specific retraining, imposing significant computational burdens and temporal constraints. To address these limitations, subsequent tuning-free paradigms [49, 88, 102, 91, 30, 98] employed pre-trained face encoders coupled with feature injection modules. These frameworks extract facial features through the encoder and subsequently integrate them into the generation process via the injection module. While these parameter-efficient methods eliminate retraining requirements and improve practical deployment capabilities, two critical limitations persist: Constraints in Multi-ID Generation: In scenarios involving multiple identities, numerous approaches struggle with the challenge of identity blending [96]. While various mitigation strategies have emerged, these methods are built upon frameworks tailored for the single-ID contexts, which inherently compromises model performance. Insufficient Facial Attribute Editability: Current methodologies lack explicit training models to disentangle identity features (e.g., facial structure and skin texture) from motion features (e.g., expression and orientation). This intrinsic entanglement within facial latent representations fundamentally restricts the flexible manipulation of facial attributes while preserving identity fidelity.
@@ 2.1 Text-to-Image Models
The field of text-to-image generation has experienced significant advancements, driven by innovative techniques such as Generative Adversarial Networks (GANs) [32, 95], auto-regressive models [68, 81, 103, 100], and diffusion models [27, 69, 97, 56, 15]. Among these techniques, diffusion models have emerged as the predominant approach, leveraging diffusion processes to iteratively transform noise into coherent images that align with textual descriptions [61, 57, 58, 79, 17, 83]. Our methodology is built upon the stable diffusion model, which is in line with current trends.
@@ 2.2 Personalized Human Image Generation
Tuning-free single-ID personalization: A vast amount of tuning-free studies have concentrated on the single-ID contexts [5, 11, 10, 31, 51, 85, 90], which typically utilize a face encoder to extract facial features from reference images and an injection module to integrate these features into the generative process. Current approaches can be systematically categorized into three distinct types based on the injection mechanism: text-based methods [82, 77, 8, 89], network-based methods [23, 102], and attention-based methods [50, 94, 19, 22, 93]. To illustrate, PhotoMaker [48] extracted facial features and merged them with the corresponding text token within the text embedding space, achieving semantic consistency though compromising identity fidelity. Conversely, InstantID [88] adapts ControlNet [101] architecture for high-fidelity generation but lacks facial editing granularity. Improving upon this, CapHuman [49] enhances facial editability via 3D Morphable Face Model [46] integration at the cost of increased computational complexity. As the pioneering attention-based method, IP-Adapter [99] introduces a novel decoupled cross-attention mechanism for facial identity integration, achieving superior fidelity while sharing InstantID’s limitation in granular facial control. Inspired by this, the $W_{+}$ adapter [47] leveraged the StyleGAN [34] encoder to extract features for expression control. However, using this encoder reduced fidelity and limited control over facial orientation. Moreover, these attention-based methods disrupted the model’s original behavior during the generative process due to the coarse-grained nature of cross-attention integration [64].
@@ 5 Conclusion
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grant 62192781, 62172326, 62137002, and in part by the Project of China Knowledge Centre for Engineering Science and Technology.